lf a line of slope 0.2 crosses the origin of x-y coordinate system and a point(x, 20) , find x

(A) 100 (B) 10 (c) 4 ( D ) 0.1 (E) 0.01

Answer by stanbon(75887)

lf a line of slope 0.2 crosses the origin of x-y coordinate system and a point(x, 20) , find x
-------------------------
Draw the picture and you will see you have two points (0,0) and (x,20)
---
slope = (20-0)/(x-0) = 20/x
----
Equation:
20/x = 0.2
Invert:
x/20 = 10/2
x/20 = 5
x = 100
==============
Cheers,
Stan H.